OpenStudy (anonymous):

PLEASE HELP: For the polynomial, list each real zero and its multiplicity. Determine whether the graph crosses or touches the x-axis at each x -intercept. f(x) = 4(x - 7)(x + 4)4

OpenStudy (jdoe0001):

\(\bf f(x) = 4(x - 7)(x + 4)^4\\ 0 =4(x - 7)(x + 4)^4\) so what values does "x" have?

OpenStudy (jdoe0001):

keep in mind that \(\bf 0 =4(x - 7)(x + 4)^4 \implies 0 =4(x - 7)(x + 4)(x + 4)(x + 4)(x + 4)\)

OpenStudy (psymon):

Oh, that was a 4th power x_x